Nancy drove 1 hour longer than Deborah at an average speed 4 kilometers
per hour faster than Deborah. Edward drove 3 hours longer than Deborah at an average speed 12 kilometers per hour faster than Deborah. If Nancy drove 80 kilometers more than Deborah, then how many more kilometers did Edward drive than Deborah?

Answer by ankor@dixie-net.com(22740)

Nancy drove 1 hour longer than Deborah at an average speed 4 kilometers
per hour faster than Deborah.
:
Edward drove 3 hours longer than Deborah at an average speed 12 kilometers per hour faster than Deborah.
:
If Nancy drove 80 kilometers more than Deborah, then how many more kilometers did Edward drive than Deborah?
:
Assign 60km/hr as Deb's speed
let t = Deb's travel time
then
60t = Deb's distance
:
64 = Nan's speed
(t+1) = Nan's time
then
64(t+1) = Nan's dist
:
72 = Ed's speed
(t+3) = Ed's time
then
72(t+3)= Ed's distance
:
If Nancy drove 80 kilometers more than Deborah,
Nan's dist = Deb's dist + 80 km
64(t+1) = 60t + 80
6tt + 64 = 60t + 80
64t - 60t = 80 - 64
4t = 16
t = 4 hrs is Deb's travel time
then
5 hrs is Nan's travel time
and
7 hrs is Ed's travel time
:
Find the distance each has traveled
4 * 60 = 240 Km
5 * 64 = 320 Km
7 * 72 = 504 km
"how many more kilometers did Edward drive than Deborah?"
504 - 240 = 264 km further